Circular No. 6488

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GX 1+4 AND GRS 1915+105 P. C. Agrawal, B. Paul, A. R. Rao, M. N. Vahia, and J. S. Yadav,

Tata Institute of Fundamental Research, Mumbai; and T. M. K. Marar, S. Seetha and K. Kasturirangan, ISRO Satellite Centre, Bangalore, report: “The x-ray pulsar GX 1+4 was observed with the Indian X-ray Astronomy Instrument onboard the IRS-P3 satellite launched from India on Mar. 21. The instrument consists of three identical, coaligned, multilayer proportional counters collimated to a 2^{o} × 2^{o} field-of-view, with an effective area of ∼ 1200 cm^{2}. GX 1+4 was observed during July 10–19 and again on Aug. 16–17. It was detected with an average countrate of ∼ 90 counts/s (equivalent to 90 mCrab) in the band 2–18 keV during this period. X-ray pulsations were weak and detectable only on July 13, Aug. 16, and 17, the pulse fraction then being 7 ± 3.5, 20 ± 5, and 6 ± 2 percent, respectively. The pulses were very broad, with a duty cycle of ∼ 80 percent on Aug. 16 and 17 and 60 percent on July 13. The upper limit on the pulsed fraction for all the other days is 4 percent. The superluminal x-ray transient GRS 1915+105 was detected with the same instrument on July 24, 25, and 27 as a bright source with an average countrate of ∼ 760 counts/s (760 mCrab) in the band 2–18 keV. Timing analysis revealed rapid intensity variations on a timescale of 0.1–0.5 s by a factor of up to 3. Strong quasiperiodic oscillations (QPOs) were detected on all days with a QPO frequency in the range 0.62–0.82 Hz; the QPOs were detected in all three detectors separately. The frequency was found to change slowly over a period of hours, and the QPO fraction varied in the range 10–20 percent. No hourly or longer-term variations were detected.”

COMET C/1996 P2 (RUSSELL-WATSON) Ephemeris continuation to IAUC 6465 (elements MPC 27882):

1996TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Oct. 14 1^{h}12.^{m}15 −35^{o}56.^{′}4 2.499 3.289 135.^{o}9 12.^{o}2 13.2
24 1 05.58 −35 21.8 2.632 3.373 131.4 12.8 13.4
Nov. 3 1 00.40 −34 23.9 2.783 3.457 125.7 13.5 13.6
13 0 56.93 −33 07.4 2.951 3.542 119.3 14.1 13.8
23 0 55.32 −31 36.9 3.134 3.628 112.4 14.6 14.1
Dec. 3 0 55.49 −29 57.0 3.329 3.713 105.3 14.8 14.3
13 0 57.32 −28 11.2 3.533 3.799  98.1 14.9 14.5
23 1 00.62 −26 22.7 3.743 3.885  90.9 14.7 14.8
1996 October 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green